# Enabling high mass accretion rates onto massive main sequence stars by outer envelope mass removal

### v1

@@ 1 Introduction
Intermediate Luminosity Optical Transients (ILOTs) are transient objects powered by mass accretion onto a main sequence or slightly evolved star. The accreted mass may come from a companion star via mass transfer or a merger process, both occurring at high rates. ILOTs typically have peak luminosities between those of classical novae, or even slightly below classical novae, and those of supernovae (e.g., Mould et al. 1990; Bond et al. 2003; Rau et al. 2007; Ofek et al. 2008; Mason et al. 2010; Kasliwal 2011; Tylenda et al. 2013; Kasliwal et al. 2012; Kaminski et al. 2018; Boian & Groh 2019; Cai et al. 2019; Jencson et al. 2019; Kashi et al. 2019; Pastorello et al. 2019; Blagorodnova et al. 2020; Banerjee et al. 2020; Howitt et al. 2020; Jones 2020; Kamiński et al. 2020, 2021; Klencki et al. 2021; Stritzinger et al. 2020a, b; Blagorodnova et al. 2021; Mobeen et al. 2021; Pastorello et al. 2021, 2023; Addison et al. 2022; Cai et al. 2022a; Wadhwa et al. 2022; Kamiński et al. 2023; Karambelkar et al. 2023; Mobeen et al. 2024; Kaminski 2024). We use the term ILOTs for all transients powered by gravitational energy, (for earlier use of this term see, e.g., Berger et al. 2009; Kashi & Soker 2016a; Muthukrishna et al. 2019). Other terms with partial (but not complete) overlap with ILOTs include luminous red novae, red novae, intermediate luminosity red transients, or gap transients (e.g., Jencson et al. 2019; Cai et al. 2022b). There is also a diversity in classifying the subclasses of these transients (e.g., Kashi & Soker 2016a; Pastorello et al. 2019; Pastorello & Fraser 2019). We consider luminous red novae to be ILOTs powered by a complete merger that leaves one stellar remnant (Kashi & Soker, 2016a), and for a common envelope evolution (CEE) because during the outburst, there is only one photosphere, that of the common envelope.
@@
High mass transfer rates from a giant star onto a main sequence companion can occur during grazing envelope evolution (GEE) and CEE. In the case of eccentric orbits, the jet-launching might be along the entire orbit or near periastron passages. Although the accreting companion might, in principle, be sub-stellar (a planet or a brown dwarf (e.g., Retter & Marom 2003; Metzger, Giannios, & Spiegel 2012; Yamazaki, Hayasaki, & Loeb 2017; Kashi et al. 2019; Gurevich, Bear, & Soker 2022; De et al. 2023; O’Connor et al. 2023), we here consider main sequence companions, as most studies do (e.g., Tylenda et al. 2011; Ivanova et al. 2013; Nandez et al. 2014; Kamiński et al. 2015; Pejcha et al. 2016a, b; Soker 2016a; Blagorodnova et al. 2017; MacLeod et al. 2017, 2018; Segev et al. 2019; Howitt et al. 2020; MacLeod & Loeb 2020; Qian et al. 2020; Schrøder et al. 2020; Blagorodnova et al. 2021; Addison et al. 2022; Zhu et al. 2023; Tylenda et al. 2024). Energy sources during the CEE might be a recombination of the ejected gas (e.g., Matsumoto & Metzger 2022), dynamical interaction of the companion with an envelope of the giant star (which causes spiraling-in of the companion), and mass accretion onto the companion.
@@
The bipolar morphology of the material that the Luminous blues variable Eta Carinae ejected in the great Eruption of the nineteenth century, the Homunculus, is compatible with a jet-powering of the Great Eruption (e.g., Soker 2001; Kashi & Soker 2010). The jet-powering scenario requires that the companion’s (secondary) mass of Eta Carinae is $M_{2}\simeq 30-80M_{\odot}$ (Kashi & Soker, 2016b), and that during the twenty-year-long Great Eruption, the companion accreted a mass of $\simeq 4M_{\odot}$ (Kashi & Soker, 2010). To power an energetic event, the companion must accrete large amounts of mass without expanding much. In a previous exploratory study, we (Bear & Soker, 2024) showed that massive main-sequence stars can avoid expansion while accreting mass at high rates if there is a mechanism to eject the outer parts of accreted mass simultaneously with mass accretion. We considered the mass removal mechanism to be jets that the star launches. The present study uses different criteria for a more realistic scenario to achieve this mass removal (Section 2). Our results strengthen the earlier finding of the mechanism to prevent stellar expansion (Section 3). In Section 4, we calculate the properties of the accretion disk, which shows it can exist in the outer zones of the envelope. We summarize in Section 6.
@@ 2 Method
We accrete mass in pulses, each comprising two parts. In the first part, we add mass at a rate of $\dot{M}_{\rm add}=1.5\times 10^{-2}M_{\odot}{~{}\rm yr}^{-1}$, and simultaneously, we deposit energy at a fraction $\eta_{\rm acc}$ of the gravitational energy that the added mass releases; we deposit this energy in the envelope’s outer $0.1$ by radius. When the star expands to a radius of
$R_{\rm add}=k_{R}R_{0}$ (1)
@@ 3 Results
The main conclusion of our simulations is that mass removal from the stellar outskirts allows massive stars to accrete up to $\simeq 10\%$ of their mass when their radius is truncated at $k_{R}=1.3$ times their initial radius, without an increase in their final stellar radius. The maximum mass is reached asymptotically. Lower stellar mass can grow by a larger factor than more massive stars. Larger maximum radii allow the star to accrete mass at a higher rate and reach higher asymptotic mass. We keep the maximum stellar radius small to maintain a deep potential well so that the accretion disk launches energetic jets. As Bear & Soker (2024) showed, the outer parts have very high entropy, such that their removal causes the envelope to shrink. Bear & Soker (2024) added and removed mass in pre-determined steps of mass addition and removal, while here we set maximum and minimum stellar radii, which we consider more realistic. When the high-entropy outer envelope regions are not removed, the star expands unstably for the accretion rates we simulate here, as Schürmann & Langer (2024) obtained in their extensive study of main sequence stars accreting at constant rates without mass removal.
@@ 4 The accretion disk
Figure 3 shows that the density in the accretion disk is larger by order of magnitude and more than the envelope density in the outer $\simeq 10\%$ of the envelope by radius. The higher disk density implies that the disk can robustly exist in the outer envelope; in our proposed scenario, the accretion disk launches jets that remove the outer envelope. We term this scenario jetted-mass-removal accretion.
@@ 5 The ejected mass
Assuming the accreted gas comes from a distance much larger than the stellar radius, with the virial theorem, the net energy that the accreted mass releases is $E_{\rm acc}=M_{\rm acc}v^{2}_{\rm es}/4$, where $v_{\rm es}$ is the escape velocity from the stellar surface. This implies that the average, by energy, velocity of the ejected material is
$\bar{v}_{\rm ej}=(2q_{\rm ej})^{-1/2}v_{\rm es}.$ (5)
@@ 6 Discussion and Summary
In Figure 4, we presented the ratio of ejected mass to net accreted mass. The jetted-mass-removal accretion involves the removal of a large fraction of the added mass when the star accretes large amounts of mass, i.e., $\gtrsim 5\%$ of its mass. This implies that the ejecta average velocity is much lower than the escape velocity from the star.
@@ Acknowledgments
A grant from the Pazy Foundation supported this research.

### v2

@@ 1 Introduction
Intermediate Luminosity Optical Transients (ILOTs) are transient objects powered by mass accretion onto a main sequence or slightly evolved star. The accreted mass may come from a companion star via mass transfer or a merger process, both occurring at high rates. ILOTs typically have peak luminosities between those of classical novae, or even slightly below classical novae, and those of supernovae (e.g., Mould et al. 1990; Bond et al. 2003; Rau et al. 2007; Ofek et al. 2008; Mason et al. 2010; Kasliwal 2011; Tylenda et al. 2013; Kasliwal et al. 2012; Kaminski et al. 2018; Boian & Groh 2019; Cai et al. 2019; Jencson et al. 2019; Kashi et al. 2019; Pastorello et al. 2019; Blagorodnova et al. 2020; Banerjee et al. 2020; Howitt et al. 2020; Jones 2020; Kamiński et al. 2020, 2021; Klencki et al. 2021; Stritzinger et al. 2020a, b; Blagorodnova et al. 2021; Mobeen et al. 2021; Pastorello et al. 2021, 2023; Addison et al. 2022; Cai et al. 2022a; Wadhwa et al. 2022; Kamiński et al. 2023; Karambelkar et al. 2023; Mobeen et al. 2024; Kaminski 2024). We use the term ILOTs for all transients powered by gravitational energy, (for earlier use of this term see, e.g., Berger et al. 2009; Kashi & Soker 2016a; Muthukrishna et al. 2019). Other terms with partial (but not complete) overlap with ILOTs include luminous red novae, red novae, intermediate luminosity red transients, or gap transients (e.g., Jencson et al. 2019; Cai et al. 2022b). There is also a diversity in classifying the subclasses of these transients (e.g., Kashi & Soker 2016a; Pastorello et al. 2019; Pastorello & Fraser 2019). We consider luminous red novae to be ILOTs powered by a complete merger that leaves one stellar remnant (Kashi & Soker, 2016a). Considering the emission processes, the common envelope evolution (CEE) might also be considered as a complete merger because there is only one photosphere, that of the common envelope. Later, the common envelope is ejected, and the core and companion emerge so from the dynamical point of view it is not a complete merger.
@@
High mass transfer rates from a giant star onto a main sequence companion can occur during grazing envelope evolution (GEE) and CEE. In the case of eccentric orbits, the jet-launching might be along the entire orbit or near periastron passages. Although the accreting companion might, in principle, be sub-stellar, i.e., a planet or a brown dwarf (e.g., Retter & Marom 2003; Metzger, Giannios, & Spiegel 2012; Yamazaki, Hayasaki, & Loeb 2017; Kashi et al. 2019; Gurevich, Bear, & Soker 2022; De et al. 2023; O’Connor et al. 2023), we here consider main sequence companions, as most studies do (e.g., Tylenda et al. 2011; Ivanova et al. 2013; Nandez et al. 2014; Kamiński et al. 2015; Pejcha et al. 2016a, b; Soker 2016a; Blagorodnova et al. 2017; MacLeod et al. 2017, 2018; Segev et al. 2019; Howitt et al. 2020; MacLeod & Loeb 2020; Qian et al. 2020; Schrøder et al. 2020; Blagorodnova et al. 2021; Addison et al. 2022; Zhu et al. 2023; Tylenda et al. 2024). Energy sources during the CEE might be a recombination of the ejected gas (e.g., Matsumoto & Metzger 2022), dynamical interaction of the companion with an envelope of the giant star (which causes spiraling-in of the companion), and mass accretion onto the companion.
@@
The bipolar morphology of the material that the Luminous blues variable Eta Carinae ejected in the great Eruption of the nineteenth century, the Homunculus, is compatible with a jet-powering of the Great Eruption (e.g., Soker 2001; Kashi & Soker 2010). The jet-powering scenario requires that the companion’s (secondary) mass of Eta Carinae is $M_{2}\simeq 30-80M_{\odot}$ (Kashi & Soker, 2016b), and that during the twenty-year-long Great Eruption, the companion accreted a mass of $\simeq 4M_{\odot}$ (Kashi & Soker, 2010). To power an energetic event, the companion must accrete large amounts of mass without expanding much. In a previous exploratory study, we (Bear & Soker, 2025) showed that massive main-sequence stars can avoid expansion while accreting mass at high rates if there is a mechanism to eject the outer parts of accreted mass simultaneously with mass accretion. We considered the mass removal mechanism to be jets that the star launches. The present study uses different criteria for a more realistic scenario to achieve this mass removal (Section 2). Our results strengthen the earlier finding of the mechanism to prevent stellar expansion (Section 3). In Section 4, we calculate the properties of the accretion disk, which shows it can exist in the outer zones of the envelope. In Section 5 we discuss some properties of the ejected mass. We summarize in Section 6.
@@ 2 Method
We accrete mass in pulses, each comprising two parts. In the first part, we add mass at a rate of $\dot{M}_{\rm add}=1.5\times 10^{-2}M_{\odot}{~{}\rm yr}^{-1}$, and simultaneously, we deposit energy at a fraction $\eta_{\rm acc}$ of the gravitational energy that the added mass releases. Specifically, we deposit energy during the mass addition part (but not the mass removal part) at a rate of $\dot{E}_{\rm add}=\eta_{\rm acc}\dot{M}_{\rm add}[GM(t)/R(t)]$, where $M(t)$ and $R(t)$ are the mass and radius of the star, respectively. We deposit this energy in the envelope’s outer $0.1$ by radius, with a uniform power per unit mass in this zone. When the star expands to a radius of
$R_{\rm add}=k_{R}R_{0},$ (1)
@@
With the specific scheme we apply the mesa stellar code cannot handle much higher mass addition rates than $\dot{M}_{\rm add}=1.5\times 10^{-2}M_{\odot}{~{}\rm yr}^{-1}$. Future one-dimensional studies will examine different mass addition and removal schemes to allow higher mass addition rates. Eventually, three-dimensional simulations should examine the maximum mass accretion rate possible.
@@ 3 Results
The main conclusion of our simulations is that mass removal from the stellar outskirts allows massive stars to accrete up to $\simeq 10\%$ of their mass when their radius is truncated at $k_{R}=1.3$ times their initial radius, without an increase in their final stellar radius. The maximum mass is reached asymptotically. Lower stellar mass can grow by a larger factor than more massive stars. Larger maximum radii allow the star to accrete mass at a higher rate and reach higher asymptotic mass. We keep the maximum stellar radius small to maintain a deep potential well so that the accretion disk launches energetic jets. As Bear & Soker (2025) showed, the outer parts have very high entropy, such that their removal causes the envelope to shrink. We demonstrate the high entropy of the accreted mass in figure 3. We present the entropy profiles as a function of mass at three times of the first pulse of the $\eta_{\rm acc}=0.1$ case that we present in Figure 1: before mass addition, at the end of mass addition, and at the end of mass removal. The outer zone of the envelope after mass addition has high entropy (red line). We remove this mass in the mass removal part. Bear & Soker (2025) added and removed mass in pre-determined steps of mass addition and removal, while here, we set maximum and minimum stellar radii, which we consider more realistic. When the high-entropy outer envelope regions are not removed, the star expands unstably for the accretion rates we simulate here, as Schürmann & Langer (2024) obtained in their extensive study of main sequence stars accreting at constant rates without mass removal.
Figure 3: Entropy profiles as a function of mass at three stages of the first accretion pulse of the $M_{\rm ZAMS}=80M_{\odot}$ model with $R_{\rm add}=1.3R_{0}$ and $\eta_{\rm acc}=0.1$. The black, red, and blue lines represent the entropy profiles of the model at $t_{\rm acc}=0$ (just before accretion starts), end of mass addition phase of the first pulse, and the end of the mass removal phase of the first pulse, respectively. Entropy is given by mesa as the specific entropy divided by $N_{\rm A}k_{\rm B}$, where $N_{\rm A}$ is Avogadro’s number and $k_{\rm B}$ is the Boltzmann constant.
@@ 4 The accretion disk
Figure 4 shows that the density in the accretion disk is larger by an order of magnitude and more than the envelope density in the outer $\simeq 10\%$ of the envelope by radius. The higher disk density implies that the disk can robustly exist in the outer envelope; in our proposed scenario, the accretion disk launches jets that remove the outer envelope. We term this scenario jetted-mass-removal accretion.
@@ 5 The ejected mass
We assume the accreted gas comes from a distance much larger than the stellar radius. The removed mass originates from large distances and is carried away by the jets back to similarly large distances; therefore, it adds no net energy to the outflow. Only the accreted mass, $M_{\rm acc}=M_{\rm add}-M_{\rm rem}$ contributed energy to the removed mass. With the virial theorem, the net energy that the accreted mass releases is $E_{\rm acc}=M_{\rm acc}v^{2}_{\rm es}/4$, where $v_{\rm es}$ is the escape velocity from the stellar surface. We assume that most of it, a fraction $\zeta$, goes to expel the removed mass, and only a small fraction $1-\zeta\ll 1$ goes to radiation. The last inequality is justified because the jet-envelope interaction occurs in a very optically thick gas (inside the photosphere). The above discussion implies that the average, by energy, terminal (i.e., at large distances) velocity of the ejected material is
$\bar{v}_{\rm ej}=\left(\frac{\zeta}{2q_{\rm ej}}\right)^{1/2}v_{\rm es}.$ (5)
@@ 6 Discussion and Summary
In Figure 5, we presented the ratio of ejected mass to net accreted mass. The jetted-mass-removal accretion involves the removal of a large fraction of the added mass when the star accretes large amounts of mass, i.e., $\gtrsim 5\%$ of its mass. This implies that the ejecta average terminal velocity (velocity at large distances relative to the stellar radius) is much lower than the escape velocity from the star.
@@ Acknowledgments
We thank an anonymous referee for helpful comments. A grant from the Pazy Foundation supported this research.